OpenStudy (anonymous):

I am in dire need of Math help. What is y-1=-1/3 (x-5) in standard form? Once solved, what is the X-intercept of the answer?

OpenStudy (texaschic101):

standard form is Ax + By = C y - 1 = -1/3(x - 5) y - 1 = -1/3x + 5/3 y = - 1/3x + 5/3 + 1 y = -1/3x + 5/3 + 3/3 y = -1/3x + 8/3 (multiply by 3) 3y = -x + 8 x + 3y = 8 <-- standard form to find the x intercept, sub in 0 for y x + 3(0) = 8 x + 0 = 8 x = 8 <-- x intercept